OpenStudy (anonymous):

Write an equation of a line whose graph is parallel to the graph of y = 3x – 10.

OpenStudy (ghazi):

it will be y=3x+d ..you can find d by substituting a pair of point at x and y

OpenStudy (anonymous):

can you give me an example ?

OpenStudy (ghazi):

like if the line would have passed through (1,2) then we have substituted (1,2) in that equation 2=3*1+d therefore d= 2-3=-1